# Results and Lessons Learned from Autonomous Driving Transportation Services in Airfield, Crowded Indoor, and Urban Environments

### 02 Mar 2024 v1

@@ Abstract
Autonomous vehicles have been actively investigated over the past few decades. Several recent works show the potential of autonomous driving transportation services in urban environments with impressive experimental results. However, these works note that autonomous vehicles are still occasionally inferior to expert drivers in complex scenarios. Furthermore, they do not focus on the possibilities of autonomous driving transportation services in other areas beyond urban environments. This paper presents the research results and lessons learned from autonomous driving transportation services in airfield, crowded indoor, and urban environments. We discuss how we address several unique challenges in these diverse environments. We also offer an overview of remaining challenges that have not received much attention but must be addressed. This paper aims to share our unique experience to support researchers who are interested in realizing the potential of autonomous vehicles in various real-world environments.
@@ I Introduction
Recently, academia and the industry have paid much attention to the great potential of autonomous vehicles. Furthermore, the public’s interest and understanding of autonomous vehicles have rapidly increased in the past few years as product-level autonomous vehicles emerge beyond the research stage and the public can now directly experience autonomous driving. Many commercial vehicles have a level-2 autonomous driving system defined by the Society of Automotive Engineers [1, 2]. Particularly, Tesla’s full self-driving (FSD) allows people to experience the autonomous driving system in daily life through their vehicles.
@@
Figure 1: Autonomous driving service area: (a) airfields: Cincinnati/Northern Kentucky Airport (2.68 km); (b) crowded indoors: Incheon Airport Terminal 1 (300 m) and Terminal 2 (550 m); and (c) urban: top: from the left, last-mile delivery in Seoul (4.67 km) and testbed in Seoul (8.21 km) and bottom: from the left, shuttle in Sejong (6.12 km), health care service in Gwangju (5.2 km), and road infra monitoring in Hwaseong (15.5 km).
Contrary to public expectations, autonomous driving systems still exhibit numerous technical challenges in complex urban environments [3, 4]. Although Tesla’s FSD beta was released at the end of 2020 and has considerably improved ever since, various incidents, including an eight-car pileup on the highway, have occurred. Since November 2022, hundreds of robotaxi fleets from Waymo and Cruise have been operating 24 h a day in San Francisco. However, these units still reveal many problems, including sudden stoppage on the road, traffic flow disruption, and blocking the path of emergency vehicles. In October 2023, the California Department of Motor Vehicles mandated the cessation of Cruise’s robotaxi service.
@@
Autonomous vehicles can be applied for diverse applications beyond urban driving. Tow tractors that transport baggage and cargo at airfields are receiving much attention because of labor shortages and rising labor costs [5]. Compact vehicles for passengers with reduced mobility have been developed for indoor environments, such as large shopping malls [6, 7]. Applying autonomous vehicles to these applications can have a significant positive effect on our society. However, it presents unique and challenging problems that are not encountered in urban environments. Furthermore, these problems have not yet been sufficiently considered in previous works.
@@
Figure 2: Developed autonomous driving platforms: (a) airfields: tow tractor; (b) crowded indoors: shuttle (AirRide); and (c) urban: top: from the left, last-mile delivery (Eligo) and on-demand taxi and bottom: from the left, health care service vehicle, road infra monitoring vehicle, and shuttle.
@@
In this paper, we present the results and lessons learned from autonomous driving transportation services in airfield, crowded indoor, and urban environments. Figs. 1 and 2 describe our service areas and vehicle platforms. In particular, we ran an autonomous tow tractor service at the Cincinnati/Northern Kentucky International Airport (CVG) airfield. We have also been operating indoor autonomous shuttle services at Incheon International Airport (ICN) arrival and departure halls. Lastly, we have been operating last-mile delivery and shuttle services in urban environments, including Seoul, Sejong, Gwangju, Hwaseong, and Jinhae in South Korea.
@@
Our work aims to share our unique experience to support researchers who are interested in deploying autonomous vehicles in diverse environments. The main contributions of our work are summarized as follows:
  • We successfully conducted autonomous driving transportation services in airfield, crowded indoor, and urban environments.
  • We present how we address diverse and unique challenges for each service.
  • We discuss the remaining challenges that have not yet received much attention but must be resolved for product-level autonomous vehicles.
@@ II Related Works
Figure 3: System architecture of the general- and mission-purpose autonomous driving.
@@
As a solution to labor shortages and rising labor costs, there is a growing focus on autonomous ground support equipment (GSE) in airfields. Wang et al. [11] introduced an efficient dispatching framework for a fleet management system of GSE. Morris et al. [5] presented an autonomous push-back tractor application, in which the push-back tractor tows aircraft from the gate to the runway and from the runway to the gate. The experimental results of these works demonstrate that autonomous GSE can improve the efficiency of operations in airfields. However, these works do not present in-depth discussions about the challenges of autonomous driving in the airfields. Unlike these works, we propose novel perception and motion planning algorithms to handle the challenges of airfield autonomous driving and provide an overview of other additional challenges that have not received much attention but must be addressed.
Another major application of autonomous driving is indoor robots. Lee et al. [12] introduced relocalization algorithms for serving robots that often fail to estimate their pose. Tsiogas et al. [13] demonstrated the autonomous pallet-moving vehicle in the factory, focusing on pallet detection and docking scenarios. Kastner et al. [14] presented human-following and -guiding applications of service robots. They proposed semantic deep reinforcement learning that robots can navigate safely in crowded environments. While these works have shown various indoor services with robots, our work discusses another service: indoor autonomous shuttle in airports. The vehicle used for this service is larger than typical indoor robots to accommodate multi-passengers and ensures their safety and comfort, which causes distinct challenges and research directions.
III Deployment of Autonomous Vehicles in Diverse Challenging Scenarios
Our autonomous driving system is a combination of a general-purpose system that is agnostic to environments and a mission-purpose system that is designed to address environment-specific problems (Fig. 3). The general-purpose system is largely divided into localization, perception, planning, and control modules. The localization module estimates the current location of the vehicles and the perception module detects surrounding objects and predicts behavior. The planning module generates the optimal trajectory based on the results of both localization and perception modules. The control module estimates the control values to follow the optimal trajectory. The mission-purpose system comprises the novel algorithms proposed in this work. The remainder of this section discusses how the proposed algorithms in the mission-purpose system handle diverse challenges for each environment in detail.
@@ III-A Airfield Environment
We operated an autonomous tow tractor for transporting passenger baggage and cargo within a 2.68 km-long airfield at the CVG. The airfield environment has unique objects, such as aircraft, tow tractors, and K-loaders, and each object has a different driving strategy. Figs. 1a and 2a show our tow tractors and the corresponding routes, respectively. We conducted a performance evaluation over three weeks, with a total distance of 2163.2 km and an operation time of 106.7 hours. Table I presents the results of the performance evaluation in the last week, where "diseng." denotes disengagement.
@@ III-A1 Sensor Fusion for Aircraft Detection
With their large sizes and irregular shapes, aircraft are one of the most difficult objects to detect in an airfield. Most previous aircraft detection algorithms assume access to top-view images [15]. However, autonomous GSE typically cannot utilize such images in real-time, and aircraft detection by using the sensors equipped with autonomous GSE has not been addressed yet. While object detection algorithms for urban environments can be applied to aircraft detection, they have not considered detecting large-size objects.
To overcome the limitations of previous algorithms, We developed a fusion-based aircraft detection algorithm that takes advantage of image-based and LiDAR-based perception algorithms [16]. We efficiently combine both algorithms by utilizing conditional random field (CRF) optimization to merge the over-segments occurring in LiDAR detection. In the first step of merging, we project each over-segment onto an image and classify the over-segments projected onto the aircraft bounding box as the primary candidates. We then construct an undirected graph G with the candidates as the vertices and the correlations between vertices as the edges. CRF optimization is used to find the optimal solution, the best label set, by minimizing the graph’s energy function. The graph’s energy function that comprised unary potentials and the clique potentials is as follows:
@@
Our experiments confirmed the benefits of fusion-based perception algorithms, achieving higher aircraft tracking and classification accuracy than other algorithms. We also observed that our algorithm is less affected by changes in aircraft size and has a higher environment-invariant ability compared to other algorithms. These experiments were based on the dataset acquired at the CVG and Dayton International Airport. Fig. 4 shows the merged over-segments of aircraft.
@@ III-A2 Probabilistic Aircraft Intention Prediction
Aircraft has top traffic priority in the airfield. When an aircraft moves to the runway from the terminal, a GSE must stop and wait. The GSE must stop when the wing walker controls the GSE movement before the aircraft moves. The key element in dealing with these situations is the estimation of both the aircraft’s pose and its intention. Previous works were, however, not appropriate for aircraft because they focused on predicting human or vehicle behavior, which considerably differs from aircraft driving characteristics [17].
To address the above challenge, we developed herein a probabilistic algorithm that can predict the aircraft’s intention based on the possible situational context information related to the aircraft movement [18]. The situational context information ($I$) is comprised of two pieces of evidence: aircraft evidence ($A$) and surrounding object evidence ($O$). An aircraft’s pose and beacon signal are defined as $A$. The information of the GSE and the ramp agent who supports the airport ramp operation are defined as $O$. Each piece of information is expressed as a Bayesian network, which is a probabilistic graphical model. Finally, the aircraft’s intention $\kappa$ is estimated using a Bayesian classifier. The joint distribution of the aircraft intention and the situational context information is presented as follows:
@@
where $\Theta$ is the training parameter set. Likelihood $P(A|\kappa,\Theta)$ and $P(O|\kappa,\Theta)$ are defined based on the statistical distribution of the training dataset. We evaluated our algorithm under various weather conditions and airport environments, thereby successfully confirming its superior performance than other algorithms. Our algorithm demonstrated performance improvements of 2.3 % in cloudy weather, 1.6 % in rainy weather, and 9.3 % in different airport environments compared to the baseline.
@@ III-A3 Efficient Occlusion-aware Risk Assessment via Simplified Reachability Quantification
In the airfield, large objects such as aircraft and K-loaders frequently generate occlusion areas. Occlusion areas are substantially dangerous for autonomous tow tractors because undetected vehicles or workers can suddenly appear on the route. However, conventional risk assessment algorithms do not take into account the occlusion areas. Several recent works have proposed occlusion-aware risk assessment algorithms, but they either fail to identify the occlusion areas in real-time or assess the occlusion risk too conservatively [19]. This can severely hinder the safety and operation efficiency of autonomous tow tractors in the airfield.
To overcome the above limitations, we introduced an efficient occlusion-aware risk assessment algorithm that can handle diverse occluded scenarios in the airfield [20]. The key component of our algorithm is the simplified reachability quantification that uses a simple distribution model over the occluded object states to estimate their risk. This model-based quantification ensures the constant time complexity of our algorithm. We evaluated our occlusion-aware risk assessment under various simulated and real-world occlusion scenarios. The experimental results demonstrated that our algorithm achieves a lower collision rate and discomfort score with lower computation complexity than that with state-of-the-art risk assessment algorithms.
@@ III-B Crowded Indoor Environment
We have been operating indoor autonomous shuttle, called AirRide, at the ICN. AirRide is required to perform efficient and comfortable driving behaviors in crowded environments. Moreover, a fail-safe system should be implemented to ensure passenger safety in various corner cases. The service routes are 300 m in the arrival hall of Terminal 1 and 550 m in the departure hall of Terminal 2. Figs. 1b and 2b show AirRide and the served routes, respectively. We conducted a performance evaluation with a total distance of 67.5 km and an operation time of 19.2 hours. Table II presents the results of the performance evaluation.
@@
Perception accuracy Localization error Range Consecutive operation time
@@
Figure 5: Object detection results in a complex urban environments. The cyan, pink, and purple boxes show a vehicle, a pedestrian, and a cyclist, respectively.
@@ III-B1 Complex and Narrow Area Collision-free Trajectory Planning
Autonomous driving in crowded indoor environments requires trajectory optimization algorithms to infer trajectories keeping a safe distance from the surrounding dynamic objects. Furthermore, the passages that autonomous vehicles should navigate in indoor environments are much narrower than standard roads. These challenges are fatal to most previous trajectory optimization algorithms that cannot consider both constraints of collision avoidance and kinodynamics in real-time. The B-spline-based trajectory optimization algorithm is a promising approach that can overcome the limitations of previous algorithms. However, applying previous B-spline-based algorithms to our indoor shuttle is not straightforward, as it has much more complex kinodynamic constraints than other types of vehicles, such as unmanned aerial vehicles.
@@ III-B2 Fail-safe System for Passenger Safety
AirRide is not for a short-term proof of concept project but for an actual indoor service. Therefore, it should be able to respond to various safety issues that may occur. AirRide is equipped with different fail-safe functions to ensure the safety of passengers and people around the vehicle. When AirRide detects a hardware or software error, it operates a basic fail-safe function to stop the vehicle and cannot be re-operated until all these errors are resolved. In addition to the basic fail-safe function, we installed service and operation perspective fail-safe functions. First, a passenger standing up can block the camera or LiDAR, leading to danger while AirRide is moving. AirRide stops when it detects the passenger’s standing-up movement through the seat pressure sensor. Second, AirRide frequently encounters passengers covering the sensors out of curiosity. We included a function that automatically detects sensor occlusion caused by foreign objects or humans and stops the vehicle. Lastly, We applied a collision avoidance system as an active safety function and external impact detection and stopping of the vehicle as a passive safety function.

### 20 Mar 2024 v2

@@ Abstract
Autonomous vehicles have been actively investigated over the past few decades. Several recent works show the potential of autonomous vehicles in urban environments with impressive experimental results. However, these works note that autonomous vehicles are still occasionally inferior to expert drivers in complex scenarios. Furthermore, they do not focus on the possibilities of autonomous driving transportation services in other areas beyond urban environments. This paper presents the research results and lessons learned from autonomous driving transportation services in airfield, crowded indoor, and urban environments [^3] Videos for our autonomous driving services are available at https://www.youtube.com/@thordrive3021. We discuss how we address several unique challenges in these diverse environments. We also offer an overview of remaining challenges that have not received much attention but must be addressed. This paper aims to share our unique experience to support researchers who are interested in exploring autonomous driving transportation services in various real-world environments.
@@ I Introduction
Recently, academia and the industry have paid much attention to the great potential of autonomous vehicles. Furthermore, the public’s interest and understanding of autonomous vehicles have rapidly increased in the past few years as product-level autonomous vehicles emerge beyond the research stage and the public can now directly experience autonomous driving. Many commercial vehicles have a level-2 autonomous driving system defined by the Society of Automotive Engineers [1]. Particularly, Tesla’s full self-driving (FSD) allows people to experience the autonomous driving system in daily life through their vehicles.
@@
Figure 1: Autonomous driving service area. (a) Airfields: Cincinnati/Northern Kentucky Airport (2.68 km). (b) Crowded indoors: Incheon Airport Terminal 1 (300 m) and Terminal 2 (550 m). (c) Urban: top: from the left, last-mile delivery in Seoul (4.67 km) and testbed in Seoul (8.21 km) and bottom: from the left, shuttle in Sejong (6.12 km), health care service in Gwangju (5.2 km), and road infra monitoring in Hwaseong (15.5 km).
Contrary to public expectations, several recent works have shown that current autonomous vehicles are occasionally inferior to expert drivers in complex urban driving scenarios. These vehicles sometimes generate awkward driving behaviors in unsignalized intersection or roundabout scenarios as they struggle to interpret interactions with surrounding vehicles. Such awkward driving behaviors can cause traffic congestion or obstruct the routes of emergency vehicles. These circumstances indicate that there are still many challenges remaining to fully realize the benefits of autonomous driving transportation services in urban environments [2, 3].
@@
Autonomous vehicles can be applied to diverse transportation services beyond urban driving. Tow tractors that transport baggage and cargo at airfields are receiving much attention because of labor shortages and rising labor costs [4]. Compact vehicles for passengers with reduced mobility have been developed for indoor environments, such as large shopping malls [5, 6]. Applying autonomous vehicles to these services can have a significant positive effect on our society. However, it presents unique and challenging problems that are not encountered in urban environments. Furthermore, these problems have not yet been sufficiently considered in previous works.
@@
Figure 2: Developed autonomous driving platforms. (a) Airfields: tow tractor. (b) Crowded indoors: shuttle (AirRide). (c) Urban: top: from the left, last-mile delivery vehicle (Eligo) and on-demand taxi, and bottom: from the left, health care service vehicle, road infra monitoring vehicle, and shuttle.
@@
In this paper, we present the results and lessons learned from autonomous driving transportation services in airfield, crowded indoor, and urban environments. Figs. 1 and 2 describe our service areas and vehicle platforms. In particular, we ran autonomous tow tractor services at Cincinnati/Northern Kentucky International Airport (CVG) airfield. We have also been operating indoor autonomous shuttle services at Incheon International Airport (ICN) arrival and departure halls. Lastly, we have been operating last-mile delivery, shuttle, and taxi services in urban environments, including Seoul, Sejong, and Jinhae in South Korea.
@@
We aim to share our unique experience to support researchers who are interested in deploying autonomous vehicles in diverse environments. The main contributions of our work are summarized as follows: 1) we successfully conducted autonomous driving transportation services in airfield, crowded indoor, and urban environments, 2) we present how we address diverse and unique challenges for each service, and 3) we discuss the remaining challenges that have not yet received much attention but must be resolved for product-level autonomous vehicles.
@@ II Related Works
Figure 3: Overall architecture of the general- and mission-purpose autonomous driving system.
@@
As a solution to labor shortages and rising labor costs, there is a growing focus on autonomous ground support equipment (GSE) in airfields. Wang et al. [10] introduced an efficient dispatching framework for a fleet management system of GSE. Morris et al. [4] presented an autonomous push-back tractor service, in which the push-back tractor tows aircraft from the gate to the runway and from the runway to the gate. The experimental results of these works demonstrate that autonomous GSE can improve the efficiency of operations in airfields. However, these works do not present in-depth discussions about the challenges of autonomous driving in the airfields. Unlike these works, we propose novel perception and motion planning algorithms to handle the challenges of airfield autonomous driving and provide several future research directions.
Another major application of autonomous driving is indoor robots. Lee et al. [11] introduced relocalization algorithms for serving robots that often fail to estimate their pose. Tsiogas et al. [12] demonstrated the autonomous pallet-moving vehicle in the factory, focusing on pallet detection and docking scenarios. Kastner et al. [13] presented human-following and -guiding service robots. They proposed a semantic deep reinforcement learning algorithm that enables robots to navigate safely in crowded environments. While these works have shown various indoor services with robots, our work discusses another service: indoor autonomous shuttles in airports. The vehicle used for this service is larger than typical indoor robots to accommodate multi-passengers and ensures their safety and comfort, which leads to distinct challenges and research directions.
III Autonomous Driving Transportation Services in Various Challenging Environments
Our autonomous driving system is a combination of a general-purpose system that is agnostic to environments and a mission-purpose system that is designed to address environment-specific problems (Fig. 3). The general-purpose system is largely divided into localization, perception, planning, and control modules. The localization module estimates the current location of the ego vehicle and the perception module detects surrounding objects and predicts their behaviors. The planning module generates the optimal trajectory based on the results of both localization and perception modules. The control module estimates the control values to follow the optimal trajectory. The mission-purpose system comprises the novel algorithms proposed in this work. In the remainder of this section, we discuss how the proposed algorithms in the mission-purpose system handle diverse challenges for each environment in detail.
@@ III-A Airfield Environment
We operated an autonomous tow tractor for transporting passenger baggage and cargo within a 2.68 km-long airfield at the CVG. The airfield environment has unique objects, such as aircraft, tow tractors, and K-loaders, and each object has a different driving strategy. Figs. 1a and 2a show our tow tractors and the corresponding routes, respectively. We conducted a performance evaluation over three weeks, with a total distance of 2,163.2 km and an operation time of 106.7 h. Table I presents the results of the performance evaluation in the third week, where "diseng." denotes disengagement.
@@ III-A1 Sensor Fusion for Aircraft Detection
With its large sizes and irregular shapes, aircraft are one of the most difficult objects to detect in airfields. Most previous aircraft detection algorithms assume access to top-view images [14]. However, autonomous GSE typically cannot utilize such images in real-time, and aircraft detection by using the sensors equipped with autonomous GSE has not been addressed yet. We also empirically observed that detection algorithms for urban driving scenarios exhibited low accuracy when applied to aircraft due to their distinct size and shape characteristics.
To overcome the limitations of previous algorithms, we developed a fusion-based aircraft detection algorithm that takes advantage of image-based and LiDAR-based detection algorithms [15]. We efficiently combine both algorithms by utilizing conditional random field (CRF) optimization to merge the over-segments occurring in LiDAR detection. In the first step of merging, we project each over-segment onto an image and classify the over-segments projected onto the aircraft bounding box as the primary candidates. We then construct an undirected graph G with the candidates as the vertices and the correlations between vertices as the edges. CRF optimization is used to find the optimal solution, the best label set, by minimizing the graph’s energy function. The graph’s energy function that comprises unary potentials and the clique potentials is as follows:
@@
Our experiments confirmed that our fusion-based algorithm outperforms other algorithms in terms of aircraft tracking and classification accuracy. We also observed that our algorithm is less affected by changes in aircraft size and has a higher environment-invariant ability compared to other algorithms. These experiments were based on the dataset acquired at the CVG and Dayton International Airport. Fig. 4 shows the merged over-segments of aircraft.
@@ III-A2 Probabilistic Aircraft Intention Prediction
Aircraft have top traffic priority in airfields. When aircraft move to the runway from the terminal, a GSE must stop and wait. The GSE also must stop when the wing walker controls the GSE movement before the aircraft moves. The key element in dealing with these situations is the estimation of both aircraft’s pose and its intention. Existing intention prediction algorithms were, however, not appropriate for aircraft because they focused on predicting human or vehicle behavior, which considerably differs from aircraft driving characteristics [16].
To address the above challenge, we developed a probabilistic algorithm that can predict the aircraft’s intention based on the possible situational context information related to the aircraft movement [17]. The situational context information ($I$) is comprised of two pieces of evidence: aircraft evidence ($A$) and surrounding object evidence ($O$). The aircraft’s pose and beacon signal are defined as $A$. The information of the GSE and the ramp agent who supports the airport ramp operation are defined as $O$. Each piece of information is expressed as a Bayesian network, which is a probabilistic graphical model. Finally, the aircraft’s intention $\kappa$ is estimated using a Bayesian classifier. The joint distribution of the aircraft intention and the situational context information is presented as follows:
@@
where $\Theta$ is the training parameter set. Likelihood $P(A|\kappa,\Theta)$ and $P(O|\kappa,\Theta)$ are defined based on the statistical distribution of the training dataset. We evaluated our algorithm under various weather conditions and airfield environments, thereby successfully confirming its superior performance compared to other algorithms. Our algorithm demonstrated performance improvements of 2.3 % in cloudy weather, 1.6 % in rainy weather, and 9.3 % in different airfield environments compared to the baseline.
@@ III-A3 Efficient Occlusion-aware Risk Assessment via Simplified Reachability Quantification
In airfields, large objects such as aircraft and K-loaders frequently generate occlusion areas. Occlusion areas are substantially dangerous for autonomous tow tractors because undetected vehicles or workers can suddenly appear on a route. However, conventional risk assessment algorithms do not take into account the occlusion areas. Several recent works have proposed occlusion-aware risk assessment algorithms, but they either fail to identify the occlusion areas in real-time or assess the occlusion risk too conservatively [18]. This can severely hinder the safety and operation efficiency of autonomous tow tractors in airfields.
To overcome the above limitations, we introduced an efficient occlusion-aware risk assessment algorithm that can handle diverse occluded scenarios in airfields [19]. The key component of our algorithm is the simplified reachability quantification that uses a simple distribution model over the occluded object states to estimate their risk. This model-based quantification ensures the constant time complexity of our algorithm. Furthermore, our algorithm can be easily integrated with any planning algorithm by leveraging the additional velocity constraints. We evaluated our occlusion-aware risk assessment under various simulated and real-world occlusion scenarios. The experimental results demonstrated that our algorithm achieves lower collision rates and discomfort scores with lower computation complexity than those of state-of-the-art risk assessment algorithms.
@@ III-B Crowded Indoor Environment
We have been operating an indoor autonomous shuttle service at ICN. The autonomous shuttle called AirRide is required to perform efficient and comfortable driving behaviors in crowded environments. Moreover, AirRide should be implemented with a fail-safe system to ensure passenger safety in various corner cases. The service routes are 300 m at the arrival hall of Terminal 1 and 550 m at the departure hall of Terminal 2. Figs. 1b and 2b show AirRide and the served routes, respectively. We conducted a performance evaluation with a total distance of 67.5 km and an operation time of 19.2 h. Table II presents the results of the performance evaluation.
@@
Perception accuracy Localization error Driving range Consecutive operation time
@@
Figure 5: Object detection results in complex urban environments. The cyan, pink, and purple boxes show a vehicle, a pedestrian, and a cyclist, respectively.
@@ III-B1 Complex and Narrow Area Collision-free Trajectory Planning
Autonomous driving in crowded indoor environments requires trajectory optimization algorithms to infer trajectories keeping safe distances from the surrounding dynamic objects. Furthermore, the passages that autonomous vehicles should navigate in indoor environments are much narrower than standard roads. These challenges are fatal to most previous trajectory optimization algorithms that cannot consider both constraints of collision avoidance and kinodynamics in real-time. B-spline-based trajectory optimization is a promising approach that can overcome the limitations of previous algorithms. However, applying previous B-spline-based algorithms to our indoor shuttle is not straightforward, as it has much more complex kinodynamic constraints than other types of vehicles, such as unmanned aerial vehicles.
@@ III-B2 Fail-safe System for Passenger Safety
AirRide is not for a short-term proof of concept project but for an actual indoor service. Therefore, it should be able to handle various safety issues that may occur. AirRide is equipped with different fail-safe functions to ensure the safety of passengers and people around the vehicle. When AirRide detects a hardware or software error, it operates a basic fail-safe function to stop and cannot be re-operated until all these errors are resolved. In addition to the basic fail-safe function, we installed service and operation perspective fail-safe functions. First, AirRide can automatically detect sensor occlusion caused by humans covering the sensors or foreign objects. In addition, it uses the seat pressure sensor to determine if passengers are standing. This is primarily due to the risk of passengers occluding the camera or LiDAR (Fig. 2b). Next, AirRide is programmed to avoid entering irrecoverable or irreversible areas where our transportation service cannot be provided, even if some people attempt to force it into these areas intentionally. The fail-safe function for detecting such areas is based on whether AirRide can generate a trajectory to return to the reference path from its current position.